OpenStudy (anonymous):

A giant pumpkin is 90% water and weighs 250 pounds. During the week long fair the pumpkin loses water from being outside in the heat. At the end of the week the pumpkin is only 80% water. How much does the pumpkin weight at the end of the week.

OpenStudy (smileyxl3):

I'm a bit rusty on this, but I think this is the answer. 250 is 90% of what = 250 / 90 = 250 / 0.9 = 277.78 277.78=100% What is 80% of 277.78 = 80% * 277.78 = 0.8 * 277.78 = 222.224 = weight at end of week

OpenStudy (anonymous):

makes sense the way you did it. I had no clue on how to start